Circular No. 7039

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998el IN ANONYMOUS GALAXY L. Germany reports the discovery by the Mount Stromlo Abell Cluster

Supernova Search Team (cf. IAUC 6639) of an apparent supernova (V = 19.6) on V and R CCD images (limiting mag V = 20.5) taken on Oct. 22 by M. MacDonald with the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1998el is located near Abell 3122 at α = 3^{h}21^{m}01^{s}.21, δ = −41^{o}33′06′′.6 (equinox 2000.0), which is 0′′.9 east of the center of its host galaxy. A nearby star has position end figures 07^{s}.77, 03′′.6. SN 1998el was not visible on similar images taken on Oct. 3 (limiting mag V = 20.0) but was confirmed by M. MacDonald with the same telescope on Oct. 23 (limiting mag V = 22.0). Finding charts are available at

http://msowww.anu.edu.au/~reiss/Abell SNSearch/.

XTE J1946+274 = 1SAX J1945.6+2721 S. Campana, Brera Astronomical Observatory; G. L. Israel and L.

Stella, Rome Astronomical Observatory; and A. Santangelo, Istituto di Fisica Cosmica e Applicazioni all’Informatica, CNR, Palermo, report on two BeppoSAX observations performed on Oct. 4.13–4.65 and 10.94–11.53 UT of the transient 15.8-s x-ray pulsar XTE J1946+274 (cf. IAUC 7014): “On Oct. 4 the BeppoSAX pointing was inaccurate, and the source straddled the edge of the MECS detectors, preventing reliable position and flux measurements. On Oct. 10–11, the source was observed on-axis at a flux of about 44 mCrab (1–10 keV); the position was determined to be α = 19^{h}45^{m}38^{s}, δ = +27^{o}21′.5 (equinox 2000.0; error radius 1′ at 95-percent confidence level). Pulsations at periods of 15.825(1) and 15.8230(1) s were detected during the first and second observations, respectively, implying a spin-up of ∼ −0.18 s/yr. The pulse profile is complex in the MECS (energy band 1–10 keV), and considerably smoother in the PDS (20–100 keV), with a pulse fraction of ∼ 26 percent. The source was clearly detected over the range 0.5–100 keV, showing a complex continuum spectrum, with an iron line at 6.6 keV (EW = 75 eV). The inferred column density of ∼ 1.6 × 10^{22} cm^{−2} translates to A(V) ∼ 8 magnitudes. The Be star proposed as the optical counterpart of the x-ray source (IAUC 7021, 7022) lies some 10′′ outside the BeppoSAX error circle. Slitless spectroscopy obtained on Sept. 16 for the brightest (V < 15) stars within the BeppoSAX error circle shows no evidence for Hα emission lines.”

1998 October 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green